# The earliest phases of CNO enrichment in galaxiesthanks: Full Tables 1 and 2 are only available at the CDS via anonymous ftp to cdsarc.u-strasbg.fr(130.79.128.5) or via http://cdsarc.u-strasbg.fr/viz-bin/cat/J/A+A/xxx/Ayyy

### v1

@@ 2.2 High-redshift systems
Recent JWST observations have captured rest-frame near-ultraviolet (UV) and optical spectra of distant galaxies exhibiting emission lines of C, N, and O from which accurate abundances relative to H have been determined (e.g., Arellano-Córdova et al., 2022; Schaerer et al., 2022; Heintz et al., 2023; Cameron et al., 2023). While most spectra obtained thus far closely resemble those of typical metal-poor star-forming galaxies in the local universe, a few luminous, low-metallicity systems display anomalous, super-solar N/O abundance ratios (e.g., Cameron et al., 2023; Marques-Chaves et al., 2024), as opposed to less extreme C/O ratios (but see D’Eugenio et al., 2023, for an extreme C-emitter at $z\simeq 12.5$). This has given rise to various speculations about the origin of such peculiar chemical features. It has been suggested that we might be either witnessing the formation of proto-globular clusters, or detecting the signatures of Wolf-Rayet or AGB stars that lose their outer layers through stellar winds, or having a glimpse into early ISM pollution from Pop III and/or very massive stars ($m_{\star}\geq 1000$ M) forming from collisions in dense clusters, or even seeing the effects of the activity of an active galactic nucleus (see Marques-Chaves et al., 2024, and references therein).
In Sect. 6, we develop GCE models tailored to reproduce the observed properties of the prototype extreme N-emitter GN-z11, resting on our benchmark model for the MW, and compare the predicted abundances with CNO abundance estimates for GN-z11 and other similar systems (Bunker et al., 2023; Cameron et al., 2023; Maiolino et al., 2023; Isobe et al., 2023; Harikane et al., 2024; Xu et al., 2024; Senchyna et al., 2024; Ji et al., 2024).
@@ 3 Chemical Evolution Models
The increasing star-to-star scatter in chemical abundance ratios at low metallicity serves as a powerful tool to understand the earliest phases of the MW chemical enrichment (Argast et al., 2004; Cescutti et al., 2015; Wehmeyer et al., 2015; Griffith et al., 2023; Vanni et al., 2023). To gain insights into the early phases of CNO element evolution, we employ a GCE model that has been extensively tested against CNO abundance data in our and external galaxies (Romano et al., 2017, 2019, 2020). While the main focus of previous work was on [Fe/H] $>-1.5$ environments, here we add a detailed treatment of the first, inhomogeneous phases of galactic chemical enrichment (Sect. 4).
@@ 3.1 The Milky Way model
Hereafter, we provide an overview of the principal features, assumptions and limitations of the adopted GCE model for the MW, which is constantly developed and improved. Further details are available in the quoted papers.
The model utilizes a multizone framework, dividing the Galactic disc into concentric annuli of 2 kpc each. It assumes that the in-situ inner halo and thick-disc components form from an initial phase of accretion of matter of primordial chemical composition (Chiappini et al., 1997, 2001). During the first $\sim$3–4 Gyr of evolution (see Spitoni et al., 2019, 2021, for the calibration of the model against asteroseismic stellar ages), an efficient burst of star formation results in rapid gas depletion, leading to the formation of the most ancient stellar populations. The thin disc forms subsequently from a second, nearly independent infall event occurring on time scales ranging from $\sim 3\>\rm Gyr$ for the inner disc, up to a Hubble time for the outer regions (Romano et al., 2000; Chiappini et al., 2001). This “inside-out” formation of the Galactic disc is required to reproduce the abundance gradients and the gas distribution along the disc (see Matteucci & Francois, 1989, and references therein).
@@ 3.2 Nucleosynthesis prescriptions
The chemical enrichment process is followed by tracking the evolution of all stable chemical species from hydrogen (H) to europium (Eu) and by considering the mass-dependent stellar evolutionary timescales, i.e., relaxing the instantaneous recycling approximation (IRA). This enables to adequately follow the evolution of chemical elements that are produced on different time-scales by stars of different initial masses and different chemical compositions.
@@
For Pop II stars, we employ stellar yields dependent on the initial mass and metallicity of the stars that allow a satisfactory fit to CNO abundance data for different MW components and early-type galaxies (Romano et al., 2019, 2020, see also Romano et al. in prep). In particular, for low- and intermediate-mass stars (LIMS) with $1\leq m_{\star}/\rm{M}_{\odot}\leq 6$, we retrieve the stellar yields for non-rotating stars from the FRUITY database (Cristallo et al., 2009, 2011, 2015). For massive stars ($13\leq m_{\star}/\rm{M}_{\odot}\leq 120$) that explode as CCSNe, we use the yields of Limongi & Chieffi (2018, their fiducial set R) computed for initial metallicities [Fe/H] = $-3$, $-2$, $-1$, and 0 and accounting for different rotational velocities (0, 150, and 300 km s-1). For type Ia SNe (SNeIa, exploding white dwarfs in binary systems), we assume the stellar yields of Iwamoto et al. (1999). SNIa primarily contribute to the bulk of iron production during later evolutionary stages and play a significant role in the decline of the $\rm[\alpha/Fe]$ ratio at metallicities higher than investigated here. The nova outburst rate is incorporated following the methodology outlined in (Romano et al., 1999, 2021, and references therein). No contribution from novae is expected before 1 Gyr has elapsed from the beginning of star formation.
3.3 High-redshift galaxies: GN-z11
As part of this study, we aim to reproduce the observed properties of GN-z11, a compact system with half-light radius $\leq$ 200 pc observed at $z\simeq 10.6$, hosting a $10^{8\sim 9.1}$ M stellar population and characterized by a low-metallicity, N-rich ISM (12+log(O/H) = $7.84^{+0.06}_{-0.05}$, N/O 3$\simeq$ 3$\times$ 3 × solar, Bunker et al., 2023; Cameron et al., 2023; Maiolino et al., 2023; Tacchella et al., 2023; Senchyna et al., 2024). While other extreme N-emitters are being found at high redshift (Marques-Chaves et al., 2024), at the moment the available data allow us to sensibly constrain only models for GN-z11. To model high-redshift galaxies, we run a simpler single-zone model. We assume that metal-free gas is quickly accreted and converted into stars within compact systems characterized by surface gas densities much higher than that of the MW. Alike for the MW, the fresh gas is accreted according to an exponentially decreasing law, ${\rm d}M_{\rm inf}/{\rm d}t\propto e^{-t/\tau}$, where $M_{\rm inf}$ is the total mass accreted and $\tau$ is the infall time scale. The gas is converted into stars according to the Schmidt-Kennicutt law and the IMF is taken from Kroupa et al. (1993), similarly to what is done for the MW (see Sect. 3.1). The adopted nucleosynthesis prescriptions for Pop II stars are those described in Sect. 3.2.
The free parameters of the model, $\tau$ and $\nu$, are fixed as to reproduce the observed properties of GN-z11. Specifically, $\tau$ is set to $\simeq 0.440$ Gyr, which corresponds to the maximum evolutionary time allowed for GN-z11, while $\nu\simeq 2.5$ Gyr-1 is adjusted as to match the observed total stellar mass, $\log(M_{\star}/\rm{M}_{\odot})=9.1^{+0.3}_{-0.4}$, at the time of the observations (Tacchella et al., 2023). In the model for GN-z11, the chemical enrichment is followed for $\approx 0.440\>\rm Gyr$, corresponding to the time span from the Big Bang to $z\simeq 10.6$. Finally, to match the observed $\rm log(O/H)+12$, we fix the dilution factor at 50 (see next section), value consistent with the results of Nandal et al. (2024b).
@@ 4 Model implementation
The role of Pop III stars is frequently overlooked in chemical evolution models. However, while at high metallicity ([Fe/H] >absent>-> -1) stars are expected to form from a thoroughly mixed medium enriched by multiple stellar populations, which effectively washes out the distinctive chemical signatures of Pop III stars, at lower metallicities the contribution from the first massive stars becomes fundamental to explain the peculiar chemical abundances of ancient, metal-poor stars that survived to the present day (Koutsouridou et al., 2023; Vanni et al., 2023, Rossi et al. 2024, submitted). On the other hand, during the initial phases of galaxy evolution, stochastic processes associated with the star formation itself become crucial in shaping the evolution and chemical enrichment history of galaxies (e.g., Caplar & Tacchella, 2019; Tacchella et al., 2020; Wang et al., 2022; Pallottini & Ferrara, 2023; Iyer et al., 2024).
To model the early chemical enrichment phases of our Galaxy we implement the GCE model described in Sect. 3 in order to account for a pre-enrichment driven by Pop III stars and a stochastic star formation process. We assume that in the initial phases of formation, the MW is built up by gas clumps with different gas densities that merge into the MW main branch. We assume that a single Pop III SN precursor is formed in each clump, with properties (mass, internal mixing level, and explosion energy) randomly selected from the Heger & Woosley (2010) model grid. The mass of Pop III stars is stochastically selected within the mass range $10\leq m_{\rm Pop\>III}/{\rm M}_{\odot}\leq 100$, while the explosion energy, $\rm E_{SN}$, is assigned by randomly choosing among four different types of supernovae, each with a specific energy range: faint ($\rm E_{51}$333 $\rm E_{51}=E_{SN}/10^{51}$= 0.3–0.6 erg), core-collapse ($\rm E_{51}=$ 1.2–1.5 erg), high-energy ($\rm E_{51}=$ 1.8–3.0 erg), and hypernovae ($\rm E_{51}=$ 5.0–10 erg). We have not included the pre-enrichment from PISNe in our analysis. In fact PISNe, with initial masses ranging from 140 to 260 M, are predicted to enrich the surrounding ISM up to $\rm[Fe/H]\approx-2$ while underproducing N, $\rm[N/Fe]<-2$ (Salvadori et al., 2019), and no stars with such low N abundances are present in our sample (see Sect. 2). The clumps are assumed to be characterized by different dilution factors ($10^{1},10^{2},10^{3},10^{4}$, see also Nandal et al., 2024a, b), defined as the ratio between the hydrogen mass and the mass ejected by Pop III SNe. This factor regulates how much the SNe ejecta is diluted by the pristine surrounding medium and is a free parameter of our model. Note that assuming different dilution factors is equivalent to consider clumps of different sizes/densities. Finally, for each dilution factor, we consider 200 realisations of the chemical enrichment.
@@ 5 Early chemical enrichment: exploiting long-lived stars in the Milky Way
In this section, we present the results of chemical evolution models for the MW. In Fig. 4 we show the predicted abundance ratios of CNO elements as a function of [Fe/H] compared with observational data (Mucciarelli et al., 2022; Yong et al., 2013; Amarsi et al., 2019). Different shades of green represent the predictions of the stochastic GCE model obtained assuming different dilution factors for the Pop III star ejecta, i.e., different sizes/densities for the star-forming clumps at the beginning of the simulation (see Sect. 4). The first thing to note is that, to reproduce the full range of [Fe/H] values covered by our stellar sample (Sect. 2.1), we need to account for different dilution factors ranging from $10^{1}$ to $10^{4}$. Specifically, the higher dilution factors allow us to reproduce the bulk of the observational data at $\rm[Fe/H]<-2$. In this case, the ejecta from the first SNe are dispersed in denser (or larger) gas knots that contain higher amounts of neutral hydrogen, resulting in a shift of [Fe/H] towards the low metallicity regime. As the dilution factor decreases, the iron released by Pop III SNe is diluted in a less dense environment, meaning the stellar clump is less compact, or in a smaller bubble, resulting in an increase in [Fe/H] abundance. Note that our assumptions on the dilution factors only affect the [Fe/H] abundances, which depend on the total amount of neutral hydrogen present in the stellar clumps, while the [X/Fe] ratios remain independent from these assumptions.
@@
As shown in Fig. 4, our stochastic GCE model successfully reproduces the scatter observed at low metallicity both for C, for N, and for O, even if it is worth noting that for $\rm[Fe/H]<-2.5$ we could only measure upper limits to the O abundances. In Fig. 4 we also compare our results with the predictions of homogeneous GCE models that do not incorporate pre-enrichment from Pop III stars. The homogeneous GCE models adopt yields for Pop II massive stars from Limongi & Chieffi (2018), computed with different initial rotational velocities ($v_{\rm rot}=$  0, 150, and 300 km s-1) and metallicities ([Fe/H] = --3, --2, --1, and 0) of the stars. Notice that for [Fe/H] <absent<-< -3, we use the yield set corresponding to the lowest [Fe/H] value. It is clearly seen that, regardless of the assumed rotation velocities, homogeneous models fail to reproduce the observed [X/Fe] abundance ratios at $\rm[Fe/H]<-2.5$. This is particularly evident for C and N and it is, at least partly, due to the fact that the yield grid of Limongi & Chieffi (2018) does not include $Z=0$ stellar models. Indeed, at higher metallicities ($\rm[Fe/H]\gtrsim-2.5$) the models reach a satisfactory agreement with the observations. There is a clear tendency to overproduce C, though. Regarding the [N/Fe] abundance ratio, we notice that models assuming massive star yields computed with different rotational velocities adequately represent stars with low N abundances ($\rm[N/Fe]<0$) across the full metallicity range. Nevertheless, even when assuming fast rotation, which boosts primary N production at low metallicities (Meynet et al., 2006; Limongi & Chieffi, 2018), these models fall short of reproducing N-enriched stars.

### v2

@@ 2.2 High-redshift systems
Recent JWST observations have captured rest-frame near-ultraviolet (UV) and optical spectra of distant galaxies exhibiting emission lines of C, N, and O from which accurate abundances relative to H have been determined (e.g., Arellano-Córdova et al., 2022; Schaerer et al., 2022; Heintz et al., 2023; Cameron et al., 2023). While most spectra obtained thus far closely resemble those of typical metal-poor star-forming galaxies in the local universe, a few luminous, low-metallicity systems display anomalous super-solar C/O or N/O abundance ratios (e.g., Cameron et al., 2023; D’Eugenio et al., 2024; Marques-Chaves et al., 2024). This has given rise to various speculations about the origin of such peculiar chemical features. It has been suggested that we might be either witnessing the formation of proto-globular clusters, or detecting the signatures of Wolf-Rayet or AGB stars that lose their outer layers through stellar winds, or having a glimpse into early ISM pollution from Pop III and/or very massive stars ($m_{\star}\geq 1000$ M) forming from collisions in dense clusters, or even seeing the effects of the activity of an active galactic nucleus (see Marques-Chaves et al., 2024, and references therein).
In Sect. 6, we develop GCE models tailored to reproduce the observed properties of the prototype extreme C- (GS-z12) and N-emitters (GN-z11), resting on our benchmark model for the MW. We compare the predicted abundances with CNO abundance estimates for GS-z12, GN-z11, and other similar systems (Bunker et al., 2023; Cameron et al., 2023; Isobe et al., 2023; Maiolino et al., 2023; D’Eugenio et al., 2024; Harikane et al., 2024; Ji et al., 2024; Senchyna et al., 2024; Xu et al., 2024).
@@ 3 Chemical evolution models
The increasing star-to-star scatter in chemical abundance ratios at low metallicity serves as a powerful tool to understand the earliest phases of the MW chemical enrichment (Argast et al., 2004; Cescutti et al., 2015; Wehmeyer et al., 2015; Griffith et al., 2023; Vanni et al., 2023). To gain insights into the early phases of CNO element evolution, we employ a GCE model that has been extensively tested against CNO abundance data of our and external galaxies (Romano et al., 2017, 2019b, 2020). While the main focus of previous work was on [Fe/H] $>-1.5$ environments, here we add a detailed treatment of the first, inhomogeneous phases of galactic chemical enrichment (Sect. 4).
@@ 3.1 The Milky Way model
Hereafter, we provide an overview of the principal features, assumptions, and limitations of the adopted GCE model for the MW, which is constantly developed and improved. Further details are available in the quoted papers.
The model utilizes a multizone framework, dividing the Galactic disc into concentric annuli of 2 kpc each. It assumes that the in-situ inner halo and thick-disc components form from an initial phase of accretion of matter of primordial chemical composition (Chiappini et al., 1997, 2001). During the first $\sim$3–4 Gyr of evolution (see Spitoni et al., 2019, 2021, for the calibration of the model against asteroseismic stellar ages), an efficient burst of star formation results in rapid gas depletion, leading to the formation of the most ancient stellar populations. The thin disc forms subsequently from a second, nearly independent infall event, on time scales ranging from $\sim 3\>\rm Gyr$ for the inner disc up to a Hubble time for the outer regions (Romano et al., 2000; Chiappini et al., 2001). This ‘inside-out’ formation of the Galactic disc is required to reproduce the abundance gradients and the gas distribution along the disc (see Matteucci & Francois, 1989, and references therein).
@@ 3.2 Nucleosynthesis prescriptions
The chemical enrichment process is followed by tracking the evolution of all stable chemical species from hydrogen (H) to europium (Eu) and by considering the mass-dependent stellar evolutionary timescales, i.e., relaxing the instantaneous recycling approximation. This enables to adequately follow the evolution of chemical elements that are produced on different time-scales by stars of different initial mass and chemical composition.
@@
For Pop II stars, we employ stellar yields dependent on the initial mass and metallicity of the stars that allow a satisfactory fit to CNO abundance data for different MW components and early-type galaxies (Romano et al., 2019b, 2020, see also Romano et al. in prep). In particular, for low- and intermediate-mass stars (LIMS) with $1\leq m_{\star}/\rm{M}_{\odot}\leq 6$, we retrieve the stellar yields for non-rotating stars from the FRUITY database (Cristallo et al., 2009, 2011, 2015). For massive stars ($13\leq m_{\star}/\rm{M}_{\odot}\leq 120$) that explode as CCSNe, we use the yields of Limongi & Chieffi (2018, their fiducial set R) computed for initial metallicities [Fe/H] = $-3$, $-2$, $-1$, and 0 and accounting for different rotational velocities (0, 150, and 300 km s-1). For type Ia SNe (SNeIa, exploding white dwarfs in binary systems), we assume the stellar yields of Iwamoto et al. (1999). SNIa primarily contribute to the bulk of iron production during later evolutionary stages and play a significant role in the decline of the $\rm[\alpha/Fe]$ ratio at metallicities higher than investigated here. The nova outburst rate is incorporated following the methodology outlined in (Romano et al., 1999, 2021, and references therein). No contribution from novae is expected before 1 Gyr has elapsed from the beginning of star formation.
3.3 High-redshift galaxies: GN-z11 and GS-z12
As part of this study, we aim to reproduce the peculiar CNO abundances of a few young systems observed at high redshift. GN-z11 is a compact object with half-light radius $\leq$ 200 pc observed at $z\simeq 10.6$, hosting a $10^{8\sim 9.1}$ M stellar population and characterized by a low-metallicity, N-rich ISM, 12+log(O/H) = $7.84^{+0.06}_{-0.05}$, N/O 4$\simeq$ 4$\times$ 4 × solar (Bunker et al., 2023; Cameron et al., 2023; Maiolino et al., 2023; Tacchella et al., 2023; Senchyna et al., 2024). Cameron et al. (2023) set a lower limit of log(C/O) $>-0.78$ to its carbon-to-oxygen ratio, yet we note that other extreme N-emitters found at high redshift have C/O ratios fairly similar to the ones of local metal-poor galaxies (Marques-Chaves et al., 2024; Schaerer et al., 2024). The galaxy GS-z12 ($z\sim 12.5$), with stellar and gaseous masses of $\sim 5\times 10^{7}$ M and $\sim 10^{7}$ M, respectively, and typical size of 100 pc has instead a carbon content unexpectedly high for its low metallicity, log(C/O) $>-0.21$ at log(O/H)+12 = $7.59\pm 0.25$ (D’Eugenio et al., 2024).
To model these high-redshift galaxies, we run single-zone GCE models. Alike for the MW, fresh gas is accreted according to an exponentially decreasing law, $\mathrm{d}M_{\mathrm{inf}}/\mathrm{d}t\propto\mathrm{e}^{-t/\tau}$, where $M_{\mathrm{inf}}$ is the total mass accreted and $\tau$ is the infall time scale. We assume that metal-free gas is quickly accreted ($\tau\sim$ 1–3 Myr) and efficiently converted into stars ($\nu\sim$ 1–10 Gyr-1) within compact systems characterized by surface gas densities much higher than that of the MW (at the end of the simulation for GN-z11, the total surface mass density is $\Sigma\simeq 7.5\times 10^{3}$ M pc-2, to be compared to values more than two orders of magnitude lower in the solar neighbourhood). The gas is converted into stars according to the Schmidt-Kennicutt law and the IMF is taken from Kroupa et al. (1993), similarly to what is done for the MW (see Sect. 3.1). The adopted nucleosynthesis prescriptions for Pop II stars are those described in Sect. 3.2. The free parameters of the model are fixed as to reproduce the observed properties of GN-z11 and GS-z12. Specifically, in the model for GN-z11 we assume the star formation history of Tacchella et al. (2023), namely, a low-level star formation episode ($\nu=$ 1 Gyr-1) lasting 40 Myr, followed by a burst ($\nu=$ 10 Gyr-1) of longer duration, 60 Myr, that forms most of the stellar mass of the system. This vigorous burst is fed by prominent infall of gas of primordial chemical composition that gives birth to a second generation of Pop III stars, overwhelming the contribution to the chemical enrichment from previous generations of Pop II stars (see Sect. 6.2). In the model for GS-z12, a shorter (10 Myr long) but similarly efficient ($\nu\sim$ 10 Gyr-1) star-formation burst is considered, resulting in a constant star formation rate of 1.6 M yr-1 and stellar mass at the time of the observations (350 Myr after the Big Bang) of log($M_{\star}/M_{\odot}$) = 7.60, in agreement with the values of $1.62^{+0.29}_{-0.24}$ M yr-1 and log($M_{\star}/M_{\odot}$) = $7.68\pm 0.19$ inferred by D’Eugenio et al. (2024). Finally, to match the low metallicity observed for GN-z11 in spite of its powerful starburst, we have to assume that a galactic outflow vents out more than a half of the stellar ejecta. This seems a reasonable request, since massive galactic outflows are actually known to be a key mechanism to regulate, or even quench, the star formation of high-redshift starbursts (e.g., Jones et al., 2019, and references therein). No galactic outflows are required for the smaller galaxy, GS-z12.
@@ 4 Model implementation
The role of Pop III stars is frequently overlooked in chemical evolution models. However, while at high metallicity ([Fe/H] >absent>-> -1) stars are expected to form from a thoroughly mixed medium enriched by multiple stellar populations, which effectively washes out the distinctive chemical signatures of Pop III stars, at lower metallicities the contribution from the first massive stars becomes fundamental to explain the peculiar chemical abundances of ancient, metal-poor stars that survived to the present day (Koutsouridou et al., 2023; Vanni et al., 2023; Rossi et al., 2024). On the other hand, during the initial phases of galaxy evolution, stochastic processes associated with the star formation itself become crucial in shaping the evolution and chemical enrichment history of galaxies (e.g., Caplar & Tacchella, 2019; Tacchella et al., 2020; Wang et al., 2022; Pallottini & Ferrara, 2023; Iyer et al., 2024).
To model the early chemical enrichment phases of our Galaxy we implement the GCE model described in Sect. 3 in order to account for a pre-enrichment driven by Pop III stars and a stochastic star formation process. We assume that in the initial phases of formation, the MW is built up by gas clumps with different gas densities that merge into the MW main branch. We assume that a single Pop III SN precursor is formed in each clump, with properties (mass, internal mixing level, and explosion energy) randomly selected from the Heger & Woosley (2010) model grid. The mass of Pop III stars is stochastically selected within the mass range $10\leq m_{\rm Pop\>III}/{\rm M}_{\odot}\leq 100$, while the explosion energy, $\rm E_{SN}$, is assigned by randomly choosing among four different types of supernovae, each with a specific energy range: faint ($\rm E_{51}$444 $\rm E_{51}=E_{SN}/10^{51}$.= 0.3–0.6 erg), core-collapse ($\rm E_{51}=$ 1.2–1.5 erg), high-energy ($\rm E_{51}=$ 1.8–3.0 erg), and hypernovae ($\rm E_{51}=$ 5.0–10 erg). We have not included the pre-enrichment from PISNe in our analysis. In fact PISNe, with initial masses ranging from 140 to 260 M, are predicted to enrich the surrounding ISM up to $\rm[Fe/H]\approx-2$ while underproducing N, $\rm[N/Fe]<-2$ (Salvadori et al., 2019), and no stars with such low N abundances are present in our sample (see Sect. 2). The clumps are assumed to be characterized by different dilution factors ($10^{1},10^{2},10^{3},10^{4}$, see also Nandal et al., 2024a, b), defined as the ratio between the hydrogen mass and the mass ejected by Pop III SNe. This factor regulates how much the SN ejecta is diluted by the pristine surrounding medium and is a free parameter of our model. Note that assuming different dilution factors is equivalent to consider clumps of different sizes/densities. Finally, for each dilution factor, we consider 200 realisations of the chemical enrichment.
@@ 5 Early chemical enrichment: exploiting long-lived stars in the Milky Way
In this section, we present the results of chemical evolution models for the MW. In Fig. 4 we show the predicted abundance ratios of CNO elements as a function of [Fe/H] compared with observational data (Mucciarelli et al., 2022; Yong et al., 2013; Amarsi et al., 2019). Different shades of green represent the predictions of the stochastic GCE model obtained assuming different dilution factors for the Pop III star ejecta, i.e., different sizes/densities for the star-forming clumps at the beginning of the simulation (see Sect. 4). The first thing to note is that, to reproduce the full range of [Fe/H] values covered by our stellar sample (Sect. 2.1), we need to account for different dilution factors555 Hydrodynamical simulations of the expansion of SN bubbles suggest that different dilution factors have to be expected in dependence of the total energy, ejecta mass, and ISM densities (e.g., Mori et al., 2002; Romano et al., 2019a, and references therein). A proper treatment of the ISM dynamics and energetic is still challenging (e.g., Kim et al., 2023) and well beyond the scope of this work. ranging from $10^{1}$ to $10^{4}$. Specifically, the higher dilution factors allow us to reproduce the bulk of the observational data at $\rm[Fe/H]<-2$. In this case, the ejecta from the first SNe are dispersed in denser (or larger) gas knots that contain higher amounts of neutral hydrogen, resulting in a shift of [Fe/H] towards the low metallicity regime. As the dilution factor decreases, the iron released by Pop III SNe is diluted in a less dense environment, meaning the stellar clump is less compact, or in a smaller bubble, resulting in an increase in [Fe/H] abundance. Note that our assumptions on the dilution factors only affect the [Fe/H] abundances, which depend on the total amount of neutral hydrogen present in the stellar clumps, while the [X/Fe] ratios remain independent from these assumptions.
@@
Figure 7: Same as Fig. 6, in the [O/Fe]–[Fe/H] space.
Refer to captionRefer to $\end{array}$
@@
Refer to captionRefer to $\end{array}$
Figure 9: Same as previous figure, but for carbon. Carbon and oxygen abundances for the unmixed RGB stars of Mucciarelli et al. (2022) have been derived as described in Sect. 2.1.1.
As shown in Fig. 4, our stochastic GCE model successfully reproduces the scatter observed at low metallicity for both C, N, and O, even if it is worth noting that for $\rm[Fe/H]<-2.5$ we could only set upper limits to the O abundances. In Fig. 4 we also compare our results with the predictions of homogeneous GCE models that do not incorporate pre-enrichment from Pop III stars. The homogeneous GCE models adopt yields for Pop II massive stars from Limongi & Chieffi (2018), computed with different initial rotational velocities ($v_{\rm rot}=$  0, 150, and 300 km s-1) and metallicities ([Fe/H] = --3, --2, --1, and 0) of the stars. Notice that for [Fe/H] <absent<-< -3, we use the yield set corresponding to the lowest [Fe/H] value. It is clearly seen that, regardless of the assumed rotation velocities, homogeneous models fail to reproduce the observed [X/Fe] abundance ratios at $\rm[Fe/H]<-2.5$. This is particularly evident for C and N and it is, at least partly, due to the fact that the yield grid of Limongi & Chieffi (2018) does not include $Z=0$ stellar models. Indeed, at higher metallicities ($\rm[Fe/H]\gtrsim-2.5$) the models reach a satisfactory agreement with the observations. There is a clear tendency to overproduce C, though. Regarding the [N/Fe] abundance ratio, we notice that models assuming massive star yields computed with different rotational velocities adequately represent stars with low N abundances ($\rm[N/Fe]<0$) across the full metallicity range. Nevertheless, even when assuming fast rotation, which boosts primary N production at low metallicities (Meynet et al., 2006; Limongi & Chieffi, 2018), these models fall short of reproducing N-enriched stars.